OpenStudy (anonymous):

2 more than the difference of a square of a number and the number gives a second number. The sum of the two numbers is 6. Find the two numbers

Directrix (directrix):

Let n be the number and y the second number. x^2 - x + 2 = y x + y = 6 -----------

Directrix (directrix):

x^2 - x + 2 = y x + y = 6 ---> y = 6 - x In the first equation, in place of y substitute 6 - x which is what y equals in the second equation x^2 - x + 2 = 6 - x Add x to both sides x^2 + 2 = 6 x ^ 2 = 4 Solve for x. There will be two values. Then, use y = 6 - x to find the y value that goes with each one of the two y values. @ilovemath334